OpenStudy (anonymous):

write the standard form of the equation and general form of the equation of the circle w/ radius r and ctr (h,k) r=10; (h,k)=(8,6). what is the standard form of the equation of this circle? simplify answer. thank you.

OpenStudy (mimi_x3):

(x-8)^2 + (x-6)^2 = 100 ( not sure if its right though xD)

OpenStudy (anonymous):

hi mimi, i thought i had the right answer just like you did but didn't know how to simplify it. i will try this...thank you so much...i needed someone else to confirm it. get back to you in a bit.

OpenStudy (anonymous):

(x-8)^2 + (x-6)^2 = 100 It's right except typo y-6 not x-6

OpenStudy (anonymous):

oh yes, that's right, i see it! thanks to you both.

OpenStudy (anonymous):

ur welcome.

OpenStudy (anonymous):

quick question..i'm looking at all the symbols and after the (x-8) (the one pointing upwards before the 2....i'm wondering where i can even find that...i don't even see it anywhere....hhmmmmm

OpenStudy (anonymous):

i'm actually using pearson's my math lab

OpenStudy (mimi_x3):

lols , its the index notation to the power of 2 \[(x-8)^{2} + (x-6)^{2} = 100\]

OpenStudy (mimi_x3):

\[(y-6)^{2}\]*

OpenStudy (anonymous):

guys sensational work

OpenStudy (mimi_x3):

sensational work ? lmao

OpenStudy (anonymous):

oh hahaha. finally figured that out! to be honest w/ you it's been a while since i've had math, i'm going a diff route now in genetics and here i am with a calculus class. lol. and i'm very sleepy doing homework at this hr. all of you are my angels right now. thank you.

OpenStudy (mimi_x3):

lols , yw xD

OpenStudy (anonymous):

oh goodness, now it's asking for the general standard for this equation of the circle. again, thank you so much mimi.

OpenStudy (anonymous):

sorry i didn't mean to ask...i actually got the answer to the general form...just surprised but i got it...logging off...thank you all.